OpenStudy (anonymous):

The slope of a line is -9 and the point (-10, -8) is on the line. Express the line in slope-intercept form.

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

-9x + -82

OpenStudy (anonymous):

y--8 =-9 ----- x--10 y+8=-9(x+10) y+8=-9x-90 y=-9x-98 make sure you check your signs

OpenStudy (anonymous):

The slope of a line is -16 and the point (10, 18) is on the line. Express the line in slope-intercept form.

OpenStudy (anonymous):

i got 16x + 178

OpenStudy (anonymous):

y-18 = -16 ---- x-10 y-18=-16(x-10) y-18=-16x+160 y=-16x+168

OpenStudy (anonymous):

y=-16x+178

OpenStudy (anonymous):

The slope of a line is 8 and the point (-6, -16) is on the line. Express the line in slope-intercept form.

OpenStudy (anonymous):

did you get that?

OpenStudy (anonymous):

i got this -16x + 32

OpenStudy (anonymous):

y+16 = 8 ----- x+6 y+16=8x+48 y=8x+32

OpenStudy (anonymous):

the slope is 8 n ot 16

OpenStudy (anonymous):

The slope of a line is -7 . The point (10, 3) is on the line. Determine the equation of this line. 10

OpenStudy (anonymous):

it should be -7 over 10

OpenStudy (anonymous):

y-3 = 7/10 --- x-10 10(y-3)=7(x-10) 10y-30=7x-70 10y=7x-40 y=7/10x-4

OpenStudy (anonymous):

The slope of a line is 1/8 . The point (8, 9) is on the line. Determine the equation of this line